40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1454     Entry time: Fri Apr 3 17:20:05 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: The 3.8kHz peak seems like the DARM RSE (not 100% sure though) 
Yoichi, Kentaro,

Last night, we took several measurements of the AO path loop TFs with various offsets/demod. phases tweaked.
The first attachment shows the AO path loop TF as a function of the offset (in counts) added to the DARM error signal.
Though it is a bit crowded plot, you can see a general tendency that the peak becomes lower in height and higher in frequency as
the DARM offset goes from negative to positive. Since the peak height also depends on the arm power and it fluctuates during the measurements,
the change is not monotonic function of the offset though.

Being suspicious of the demodulation phase of the DARM error signal (AS166), we scanned it (see the second attachement).
But there is no significant change.
Note that the phase of the TF is 180 degrees different from the first attachment. This is because I changed the measurement point of the returning signal
on the CM board from TP2A to OUT2 to see POX_1I signal as well. These points should give the same signal for PO_DC except for the sign.

We also took the AO path TFs by changing the MICH offset (the third attachment). Again, there is no big change.

With the CARM locked with the PO_DC signal, we took the transfer function from the AO path actuation signal to the response of the POX_1I (4th attachment).
There is a huge 3.8kHz peak.

Finally, we measured the DARM response by exciting the ETMs differentially (the PDF attachment).
The shape of the 3.8kHz resonance looks like the DARM RSE peak.

It is speculated that somehow the DARM RSE resonance is coupled into the CARM loop. Don't know how though.
I'm now working on an Optickle simulation to get an insight into this issue.
Attachment 1: AO-TF-DARM-OFFSET.png  207 kB  | Hide | Hide all
AO-TF-DARM-OFFSET.png
Attachment 2: AO-TF-DARM-DEMOD-PHASE.png  146 kB  | Hide | Hide all
AO-TF-DARM-DEMOD-PHASE.png
Attachment 3: AO-TF-MICH-OFFSET.png  141 kB  | Hide | Hide all
AO-TF-MICH-OFFSET.png
Attachment 4: POX_1I.png  173 kB  | Hide | Hide all
POX_1I.png
Attachment 5: DARM-Loop.pdf  8 kB  | Hide | Hide all
DARM-Loop.pdf
ELOG V3.1.3-